# Draft Model Knows When to Stop: A Self-Verification Length Policy for Speculative Decoding

### v1

@@ Abstract
Speculative Decoding (SD) has become an important technique in accelerating the inference speed of large language models. Conventional SD methods employ a fixed draft length, which ignores the token generation difficulty across tasks. Consequently, in this paper, we address such an issue and introduce SVIP - a difficulty-aware dynamic draft length policy for speculative decoding systems. Based on a theoretical lower bound of draft token acceptance rate and its inference-time approximation, SVIP adaptively determines the lengths of draft sequences based on the entropy of each draft token distribution. Experimental results on mainstream SD benchmarks and frameworks demonstrate the superior performance of SVIP, achieving up to 20% walltime speedup on SpecBench over baseline SD methods and 60% speedup on MT-Bench for long-form generation of up to 8K tokens. Moreover, SVIP is totally training-free and compatible with any existing SD methods that generate draft tokens autoregressively. Experimental results also show that SVIP yields consistent walltime improvement on top of GliDe & CaPE and EAGLE-2.
@@
A Self-Verification Length Policy for Speculative Decoding
Ziyin Zhang1,2thanks: Work done during their internship at Tencent AI Lab. Jiahao Xu2 Tian Liang2 Xingyu Chen1,2∗ Zhiwei He1,2∗ Rui Wang1thanks: Corresponding author.Zhaopeng Tu2 1Shanghai Jiao Tong University 2Tencent AI Lab 1{daenerystargaryen,galaxychen,zwhe.cs,wangrui12}@sjtu.edu.cn 2{jettexu,ttianliang,zptu}@tencent.com
@@ 1 Introduction
Speculative decoding (Leviathan et al., 2023; Chen et al., 2023) is a novel technique that markedly enhances the generation wall-time of large language models (LLMs). This approach employs a small and efficient amateur model to draft sequences, while concurrently utilizing a larger and more powerful expert model to verify the drafts. By avoiding the autoregressive generation of each token through the target LLM, speculative decoding achieves improved efficiency while preserving the quality of the output.
Figure 1: The “difficulty” of tokens varies in a sequence, resulting in different numbers of accepted draft tokens at different positions.
Many variants of speculative decoding have been proposed. A line of work focuses on developing a stronger and faster draft model (Li et al., 2024b; Elhoushi et al., 2024; Du et al., 2024). Another line of work contributes to maximizing the acceptance probability of draft tokens (Sun et al., 2023; Li et al., 2024a; Lu et al., 2024). In general, they all tend to maximize the alignment between the draft and target model to further maximize the system acceptance rate.
Though successful, most of these works limit their settings to a fixed draft length, where the draft model always generates a fixed number of tokens (e.g. 4 or 5) in each iteration. Such a setting ignores the fact that some tokens - such as stop words or civilities - in the generation may be easy for the draft model to predict, while others - such as knowledge-intensive or reasoning-intensive tokens - can be much harder, as shown in Figure 1.
Figure 2: The correlation between draft model entropy and draft token acceptance probability (top) and lengths of accepted draft seqeunces (bottom).
To address this issue, in this work we introduce SVIP - Self-VerIfication length Policy, a simple, plug-and-play dynamic draft length policy for speculative decoding systems, which enhances the wall time speedup of these systems by adaptively allowing for longer draft sequences for “simple” tokens (top of Figure 1) and terminating the drafting process early upon encountering “hard” tokens (bottom of Figure 1).
Specifically, we first analyze the acceptance rate of speculative decoding systems and derive a lower bound based on the system’s entropy information. Further empirical analysis suggests that such a bound can be approximated by the entropy of the draft model only, which is naturally available in the drafting process of any auto-regressive draft models. Consequently, we develop SVIP  which controls the length of draft sequences dynamically by determining whether to continue drafting or start verification upon the generation of each draft token.
With extensive experiments across multiple model sizes and generation lengths, we demonstrate the superior performance of SVIP. It yields more than 20% of improvements over vanilla speculative decoding for Qwen2.5 14B and LLaMA-3 70B on SpecBench (Xia et al., 2024), and more than 60% of improvements for Pythia 6.9B on MT-Bench (Zheng et al., 2023) when generating long-form responses of up to 8K tokens.
Moreover, since our method is lightweight and training-free, it is extremely flexible and can be adapted to any speculative decoding system with an auto-regressive draft model. As examples, we apply SVIP on top of two state-of-the-art speculative decoding methods: GliDe with a CaPE (Du et al., 2024) and EAGLE-2 (Li et al., 2024a), and confirm that it brings consistent improvements. Our code is available at https://github.com/Geralt-Targaryen/SVIP.
@@
    We derive a low bound of speculative decoding systems, where the acceptance rate of the draft model could be modeled by its entropy only.
@@
    Based on this lower bound, we further develop an entropy-based dynamic draft length policy for speculative decoding systems, which is extremely flexible and can be adapted to any auto-regressive draft model.
@@
    Experimental results demonstrate the superior performance of SVIP over baseline methods, with up to 20% average speedup on SpecBench, 60% in long-form generation, and consistent improvement over state-of-the-art speculative decoding frameworks such as GliDe & CaPE and EAGLE-2.
2 Method
Figure 3: Distribution histograms of the entropy ratio $H_{q,p}/H_{q}$. For most tokens, this ratio falls into a narrow range, indicating that the cross entropy $H_{q,p}$ can be approximated by a constant multiplication of the draft entropy $H_{q}$.
The overall objective of SVIP is to dynamically adapt draft length on-the-fly, stopping early if the current draft token’s acceptance probability is low and otherwise continuing drafting. To introduce our method, we first provide the background on speculative decoding in Section 2.1. Then, based on the key observation that the acceptance probability of a draft token depends on the target model confidence - which is unavailable in the drafting phase, we derive a theoretical lower bound for the acceptance rate based on the draft model’s entropy and the cross-entropy between the target and draft models in Section 2.2. Finally, based on empirical analysis of the target and draft distributions of more than 100K tokens across three model families, we approximate this lower bound using only the entropy information of the draft model in Section 2.3, making it viable in actual inference.
2.1 Preliminaries on Speculative Decoding
Suppose we have two LLMs $p$ and $q$, where $p$ is the larger (target) model, and $q$ is the smaller (draft) model. Given an input sequence $x_{\leqslant t}$ of length $t$, and a draft length $\gamma$, the draft model first samples $\gamma$ tokens $x_{t+1},\cdots,x_{t+\gamma}$ autoregressive, which are verified by the target models in parallel to acquire the confidences $p(x_{t+1}),\cdots,p(x_{t+\gamma})$ [^1] The short hand $p(x_{n})$ is used to denote the conditional probability $p(x_{n}|x_{<n})$ when there is no ambiguity. Throughout the work we use subscripts to indicate token indices in a sequence (e.g. $x_{n}$ for $n$-th token), and superscripts to indicate element indices in a vector (e.g. $x_{n}^{i}$ for $i$-th element in $x_{n}$)..
[…]
The complete algorithms for speculative decoding are given in Appendix A.
Figure 4: Comparison between the actual acceptance probability from Equation (2), the acceptance probability lower bound from Equation (5), and the estimated lower bound after approximating the cross entropy $H_{q,p}$ with a constant multiplication of $H_{q}$. Each position on the x-axis corresponds to a token, which has been sorted according to the actual acceptance probability.
2.2 Theoretical Lower Bound of Acceptance Rate
@@
From Section 2.1, it’s easy to derive that given an input sequence $x_{<t}$ and a draft token $x_{t}$, its acceptance probability is $\min\left(1,\;\frac{p(x_{t})}{q(x_{t})}\right)$. Let $\beta$ denote the expected acceptance probability over the distribution of $x_{t}$, and it follows that
$\beta$ $=\sum_{x}q(x)\cdot\min\left(1,\frac{p(x)}{q(x)}\right)$ (1)
$=\sum_{x}\min\left(p(x),q(x)\right).$ (2)
Chen et al. (2023) has proven that $\beta$ is related to the total variational distance (TVD) between $p$ and $q$:
$\beta=1-\mathrm{TVD}(p,q).$ (3)
According to Pinsker’s inequality - which relates TVD to Kullback-Leibler divergence - we then have
$\beta$ $\geqslant 1-\sqrt{\frac{1}{2}\mathrm{KL}(q||p)}$ (4)
$=1-\sqrt{\frac{1}{2}\sum_{x}q(x)\log\frac{q(x)}{p(x)}}$
$=1-\sqrt{\frac{1}{2}H_{q,p}-\frac{1}{2}H_{q}},$ (5)
where $H_{q,p}$ is the cross entropy between $q$ and $p$, and $H_{q}$ is the entropy of $q$.
Equation (5) provides a theoretical lower bound for the acceptance rate of a draft token using the entropy information of the speculative decoding system [^2] An alternative way to construct the lower bound is discussed in Appendix B.. To provide an intuitive motivation for using entropy to construct the lower bound, we plot the relation between draft model entropy and draft token acceptance probability as well as accepted draft seqeunce lengths in Figure 2, which are collected from more than 1M tokens generated by three different target models with temperature set to 1 and maximum draft length set to 40. The first subfigure indicates a strong negative correlation between draft model entropy and draft token acceptance probability, while the second subfigure shows that when the entropy is low, dozens of consecutive draft tokens could be accepted, which highlights the drawback of setting draft length to a constant, small value, as is the common practice in speculative decoding literature. Another version of Figure 2 in the greedy setting is given in Appendix D.
2.3 Empirical Estimation of Acceptance Rate
So far in this section, we have been assuming access to the target distribution $p(x_{t})$ of the next token, which is unavailable in the drafting phase at inference time. Thus, to apply Equation (5) for actual acceptance rate estimation, we must approximate the cross entropy $H_{q,p}$ with information from only the draft model.
@@ 2.3 Empirical Estimation of Acceptance Rate
28:$x_{\leqslant n}$
To tackle this issue, we first plot the relationship between $H_{q,p}$ and $H_{q}$ in Figure 3. For all three model families, the entropy ratio $H_{q,p}/H_{q}$ are concentrated in a narrow range between 1.0 and 1.3. Thus, we choose to approximate $H_{q,p}$ with a constant multiplication of $H_{q}$. Plugging it into Equation (5), we now have
$\beta$ $\geqslant 1-\sqrt{cH_{q}},$ (6)
where $c$ is a constant controlling the approximation ratio between $H_{q,p}$ and $H_{q}$. In Figure 4, we visualize the values derived from Equation (2), (5), and (6).
With Equation (6) providing a way to estimate acceptance probability using only the draft model’s entropy, we can now adapt the draft length on-the-fly. After generating each draft token, we compute the estimated acceptance probability lower bound, and stop the draft process if it’s lower than a certain threshold $h$. We note that since both $c$ and $h$ are constant hyperparameters, we can remove $\sqrt{c}$ from Equation 6 and absorb it into the threshold $h$.
We formalize SVIP in Algorithm 1. The details of the methods Verify and Correct are given in Appendix A, for which different versions are available for sampling (Algorithm 2, 4) and greedy decoding (Algorithm 3, 5).
@@ 3 Experiments
3.1 Experiments on SpecBench
3.1.1 Settings
Methods MT-Bench Trans. Sum. QA Math RAG Avg.
Pythia (6.9B, ${}_{\text{(6.9B, 160M)}}$ Const. 1.45 1.47 1.24 1.43 1.52 1.42 1.42
Heuristics 1.51 1.58 1.34 1.58 1.64 1.51 1.53
SVIP 1.63 1.62 1.45 1.67 1.72 1.66 1.63(+14.8%)
Qwen2.5 (14B, ${}_{\text{(14B, 0.5B)}}$ Const. 1.08 0.87 1.11 0.92 1.43 0.99 1.07
Heuristics 1.10 0.91 1.10 0.92 1.34 1.03 1.07
SVIP 1.33 1.12 1.37 1.14 1.57 1.23 1.29(+20.6%)
LLaMA-3 (70B, ${}_{\text{(70B, 8B)}}$ Const. 2.04 2.48 2.56 2.34 2.32 2.28 2.34
Heuristics 2.30 3.13 3.33 2.61 2.52 2.63 2.76
SVIP 2.31 3.04 3.48 2.63 2.89 2.59 2.83(+20.9%)
Table 1: Speedup over target-model-only autoregressive decoding on SpecBench.
We validate the effectiveness of SVIP on Spec-Bench (Xia et al., 2024) using three distinct target models: Pythia 6.9B (Biderman et al., 2023), Qwen2.5 14B (Yang et al., 2024), and LLaMA-3 70B (Dubey et al., 2024), with Pythia-160M, Qwen2.5 0.5B, and LLaMA-3 8B as the draft models respectively.
@@ 3.1.1 Settings
As baselines, we consider two simple policies for draft length: 1) a constant draft length of 5, which is commonly used in the literature, and 2) the heuristics implemented in Hugging Face Transformers library (Wolf et al., 2019), where the draft length for the next draft iteration is increased by 2 if all draft tokens in the current iteration are accepted, and otherwise decreased by 1.
We set the sampling temperature to 0 on SpecBench (the alternatives are discussed in Appendix C). For each model, the entropy threshold $t$ in SVIP is chosen from $\{0.2,0.3,0.4,0.5\}$ based on performance on 8 samples held out from MT-Bench (Zheng et al., 2023), which are 0.4 for Pythia, and 0.3 for Qwen2.5 and LLaMA-3. All experiments with Pythia and Qwen are conducted on a single 40GB A100, while experiments with LLaMA are conducted on 5 40GB A100s. To mitigate the impact of system performance varitations, we repeat all experiments with Pythia and Qwen for three times (using different random seeds when they are used) and report the average speedup over target-model-only autoregressive decoding. Also, since the memory consumption of verifying $n$ draft tokens is quadratic in $n$, we limit the maximum draft length to 40 in both heuristics and SVIP scenarios, beyond which we start to encounter out-of-memory issues.
3.1.2 Results
The results on SpecBench are shown in Table 1. Compared with the constant approach, SVIP yields an average speedup of 15% for Pythia and 20% for Qwen and LLaMA. Compared with the heuristic approach, SVIP also gives a consistent improvement on all domains for Pythia and Qwen, and outperforms the latter on 4 out of 6 domains for LLaMA.
Figure 5: The average generated draft length, accepted draft length, and acceptance rate of Qwen2.5 (top) and LLaMA-3 (bottom) on SpecBench. Compared with the two baselines, SVIP leads to a shorter draft length and a much higher acceptance rate.
In Figure 5, we plot the average draft length and accepted draft length of Qwen and LLaMA (the results for Pythia are similar, and are given in Appendix C). From the figure, we observe that by terminating the draft process when the draft model entropy is high, SVIP leads to shorter draft lengths and a much higher acceptance rate (close to or more than 80% on all domains for Qwen, and more than 90% for LLaMA). Notably, the acceptance rate of LLaMA-3 even reaches 99% on the summarization domain, which contributes to the highest speedup (3.48) in Table 1.
3.2 Long-form Generation

### v2

# Draft Model Knows When to Stop: Self-Verification Speculative Decoding for Long-Form Generation

@@ Abstract
Conventional speculative decoding (SD) methods utilize a predefined length policy for proposing drafts, which implies the premise that the target model smoothly accepts the proposed draft tokens. However, reality deviates from this assumption: the oracle draft length varies significantly, and the fixed-length policy hardly satisfies such a requirement. Moreover, such discrepancy is further exacerbated in scenarios involving complex reasoning and long-form generation, particularly under test-time scaling for reasoning-specialized models. Through both theoretical and empirical estimation, we establish that the discrepancy between the draft and target models can be approximated by the draft model’s prediction entropy: a high entropy indicates a low acceptance rate of draft tokens, and vice versa. Based on this insight, we propose SVIP: Self-Verification Length Policy for Long-Context Speculative Decoding, which is a training-free dynamic length policy for speculative decoding systems that adaptively determines the lengths of draft sequences by referring to the draft entropy. Experimental results on mainstream SD benchmarks as well as reasoning-heavy benchmarks demonstrate the superior performance of SVIP, achieving up to 17% speedup on MT-Bench at 8K context compared with fixed draft lengths, and 22% speedup for QwQ in long-form reasoning.
@@
Self-Verification Speculative Decoding for Long-Form Generation
Ziyin Zhang1,2,thanks: Work done during their internship at Tencent. Jiahao Xu2 Tian Liang2 Xingyu Chen1,211footnotemark: 1 Zhiwei He1,211footnotemark: 1 Rui Wang1thanks: Corresponding authors.Zhaopeng Tu [^2] footnotemark: 2 1Shanghai Jiao Tong University     2Tencent 1{daenerystargaryen,galaxychen,zwhe.cs,wangrui12}@sjtu.edu.cn 2{jettexu,ttianliang,zptu}@tencent.com
Figure 1: The variance of oracle draft length drastically increases with context length. MTB (MT-Bench): a conventional benchmark for SD systems. AIME: an extremely difficult mathematical testing set for advanced reasoning models.
@@ 1 Introduction
Speculative decoding (Leviathan et al., 2023; Chen et al., 2023) is a novel technique that markedly enhances the generation wall-time of large language models (LLMs). This approach employs a small and efficient draft model to draft sequences, while concurrently utilizing a larger and more powerful expert model to verify the drafts. By avoiding the autoregressive generation of each token through the target LLM, speculative decoding achieves improved efficiency while preserving the quality of the output. This technique is particularly beneficial in the context of inference-time scaling, where LLMs generally generate long-form text.
The majority of research on speculative decoding focuses on improving the acceptance rate of the draft sequences (Sun et al., 2023; Li et al., 2024b; Elhoushi et al., 2024; Du et al., 2024; Li et al., 2024a; Lu et al., 2024). However, they limit their settings to a fixed draft length (e.g. less than 5 tokens), which we find is sub-optimal in the scenario of long-form text generation.
Figure 2: Overview of SVIP: the draft model proceeds the generation process (marked by green) until it encounters a token for which it has low confidence (marked by red), signaled by high entropy, at which point the draft model would cease generation and send the draft spans to the target model for verification.
Specifically, as discussed in Section 2.1, we investigate the token rejection phenomenon in speculative decoding (SD) systems. Our findings reveal that the oracle draft length varies considerably for long-form generation, as shown in fig.˜1. Furthermore, heuristic methods, such as predicting the draft length, are impractically challenging because our further investigation reveals that token rejection occurs unexpectedly. Moreover, our closer investigation indicates a strong correlation between token rejection and the draft model’s prediction entropy at that moment.
Inspired by such a correlation, we analyze the acceptance rate and propose our SVIP: Self-Verification Length Policy in Section 2.2. Specifically, we derive a lower bound for the acceptance rate based on the entropy information from the draft model. Notably, SVIP not only approximates this lower bound but also dynamically adjusts the length of draft sequences by determining whether to continue drafting or initiate verification after each token generation, as shown in Figure 2. By optimizing draft sequence lengths, SVIP enhances SD systems’ efficiency. Importantly, our method is entirely training-free and thus can be seamlessly integrated with any SD decoding algorithm, making it broadly applicable and efficient.
With extensive experiments across multiple model sizes and evaluation benchmarks, we demonstrate the superior performance of SVIP  in long-context generation. Compared with fixed-length draft policies, it yields up to 17% speedup on MT-Bench (Zheng et al., 2023) and 22% on AIME. As a training-free length policy, SVIP is also extremely flexible and compatible with state-of-the-art speculative decoding systems such as EAGLE-2 (Li et al., 2024a), achieving an additional 13% speed improvement.
@@
    We provide an in-depth analysis of the disagreement between draft model and target model in speculative decoding systems, highlighting the underperformance of fixed-length draft length policies.
@@
    Based on this analysis, we derive a low bound of speculative decoding systems, where the acceptance rate of the draft model could be modeled by its entropy only. We further develop SVIP, an entropy-based dynamic draft length policy for speculative decoding systems, which is extremely flexible and can be adapted to any auto-regressive draft model.
@@
    Experimental results demonstrate the superior performance of SVIP over baseline draft length policies on both conventional long-form generation and reasoning-heavy benchmarks.
2 Draft Model Knows When to Stop
In this section, we first examine the behavior of draft models at the rejection phenomenon, and analyze the oracle lengths for SD systems. Then, we theoretically derive SVIP, which approximates the draft token acceptance rate using the draft model’s own prediction entropy.
2.1 Investigation of Rejection
Speculative decoding enhances the efficiency of large language model (LLM) inference by assuming that draft tokens are accepted by the LLM, thus avoiding autoregressive generation. Should the target model exhibit a tendency to reject tokens, the overall performance of the system may experience considerable degradation. Consequently, empirically investigating the rejection phenomenon is our primary interest.
Specifically, we analyze the distribution characteristics of rejected tokens across two scenarios:
  • AIME (2022-2024): a challenging math reasoning dataset, using greedy decoding with QwQ-32B-Preview (Team, 2024) and a 1.5B draft model;
  • MT-Bench: a conversational and instruction-following benchmark, using sampling decoding with the Qwen2.5 family (Yang et al., 2024).
(a) KL Div. ($\downarrow$)
(b) Sorted vocab. distribution
Figure 3: Agreement scores and sorted vocabulary log probability at the rejection phenomenon. The x-axis of the Figure˜3(a) represents rejected tokens and the four tokens before them.
Rejection occurs out of the blue
A natural question is: How does the token rejection occur? Is there any symptom before it happens? We investigate the rejected tokens with the following metrics:
  • KL divergence for vocabulary distribution difference.
  • Vocabulary distribution of accepted and rejected tokens.
We quantify the occurrence of the token rejection phenomenon along with its corresponding prefix tokens, as illustrated in Figures˜3(a) and 3(b). It shows that KL metrics experienced a sudden and substantial surge in the position of the rejected token, and the output vocabulary distribution at rejected tokens differs significantly from previously correctly drafted tokens. Such a phenomenon is expected: KL divergence indicates a system discrepancy with rejection, where the ground-truth tokens are inherently difficult for the draft model to predict. We also quantify the entropy of the draft model at the acceptance and rejection positions in Table 1. It shows that the draft models suffer from a severely high entropy at rejection, indicating extreme difficulty in modeling the corresponding tokens.
Dataset Vocab Entropy
Accepted. Rejected
AIME 0.25 1.30
MT-Bench 2.18 3.99
Table 1: Vocabulary entropy of the draft model at accepted and rejected draft tokens.
2.2 Our Method: SVIP
Since the rejection phenomenon occurs suddenly and the draft model suffers from high entropy at the rejection position, can we detect the rejection with KL divergence or draft model’s entropy? To achieve this, we seek a theoretical understanding of the rejection phenomenon.
Lower Bound of Acceptance
@@ 2.2 Our Method: SVIP
Since rejection denotes the sudden decrease of acceptance rate, we investigate the theoretic acceptance rate of the SD systems. Specifically, given a target model $p$, a draft model $q$, an input sequence $x_{<t}$, and a draft token $x_{t}$, it’s easy to derive that $x_{t}$’s acceptance probability is $\min\left(1,\;\frac{p(x_{t})}{q(x_{t})}\right)$ (see Appendix A). Let $\beta$ denote the expected acceptance probability over the distribution of $x_{t}$, and it follows that:
$\beta$ $=\sum_{x}q(x)\cdot\min\left(1,\frac{p(x)}{q(x)}\right)$
$=\sum_{x}\min\left(p(x),q(x)\right),$ (1)
where $p$ and $q$ denote the target and draft model respectively. Chen et al. (2023) has proven that $\beta$ is related to the total variational distance (TVD) between $p$ and $q$. Start from this, we utilize Pinsker’s inequality in Equation˜3 and yield the following bound in Equation˜4:
$\beta$ $=1-\mathrm{TVD}(p,q)$ (2)
$\geqslant 1-\sqrt{\frac{1}{2}\mathbb{KL}(q||p)}$ (3)
$=1-\sqrt{\frac{1}{2}H_{q,p}-\frac{1}{2}H_{q}}$ (4)
where $H_{q,p}$ is the cross entropy between $q$ and $p$, and $H_{q}$ is the entropy of $q$. We denote the above bound as the oracle bound. Utilizing this bound for acceptance prediction is infeasible since it requires instantaneous access to the target model for cross entropy $H_{q,p}$, which is infeasible during the drafting phase.
Approximating the Oracle Bound with Draft Distribution
Can we approximate the cross-entropy $H_{q,p}$ between the draft model’s distribution $q$ and the target distribution $p$ using only $q$? We propose using the draft model’s entropy $H_{q}$ as a proxy, approximating $H_{q,p}$ as $\gamma H_{q}$, where $\gamma=H_{q,p}/H_{q}$ is a random variable capturing the ratio between $H_{q,p}$ and $H_{q}$. This leads to a bound on the acceptance rate $\beta$:
$\beta\geq 1-\sqrt{\frac{1}{2}(\gamma-1)H_{q}}.$ (5)
To make this bound practical, we approximate $\gamma$ with a constant $c$, yielding the approximation bound:
$\beta\geq 1-\sqrt{\frac{1}{2}(\gamma-1)H_{q}}\approx 1-\sqrt{cH_{q}}.$ (6)
This bound holds when $1-\sqrt{\frac{1}{2}(\gamma-1)H_{q}}\geq 1-\sqrt{cH_{q}}$, ensuring the approximation is conservative. Thus, the approximation bound $1-\sqrt{cH_{q}}$ lower-bounds the true acceptance rate $\beta$.
Detecting Rejection with Draft Entropy
With Equation˜6 providing a way to estimate the acceptance probability using only the draft model’s entropy, we introduce SVIP, which dynamically adapts the draft length. After generating each draft token, we compute the approximation bound and halt drafting if it falls below a threshold $\hat{h}$, i.e., if $1-\sqrt{cH_{q}}<\hat{h}$. Since $c$ and $\hat{h}$ are constant hyperparameters, we simplify the criterion by defining a new threshold $h=(1-\hat{h})/\sqrt{c}$, absorbing $\sqrt{c}$ into $h$. This reduces the stopping condition to $\sqrt{H_{q}}>h$. Formally, given a prefix $x_{<t}$ of t1t-1 tokens, the stopping criterion is:
$\sqrt{H_{q}(x_{<t})}>h,$ (7)
where $H_{q}(x_{<t})$ is the entropy of the draft distribution conditioned on $x_{<t}$. This ensures drafting stops when the estimated acceptance probability is too low, optimizing efficiency while maintaining reliability. We formalize SVIP in Algorithm 1. The details of the methods Verify and Correct are given in Appendix A, for which different versions are available for sampling (Algorithm 2, 4) and greedy decoding (Algorithm 3, 5).
@@
[…]
Figure 4: Comparison between the actual acceptance probability in Equation˜1, the acceptance probability lower bound in Equation˜4, and the estimated lower bound in Equation˜6. Each position on the x-axis corresponds to a token, which has been sorted according to the actual acceptance probability.
Figure 5: The SD system speedup on MT-Bench using Qwen2.5-14B (top) and Qwen2.5-32B (bottom) as targets and three different smaller models as drafts.
Justifying the Approximation Bound
The tightness of the approximation depends on how well $c$ captures the behavior of the random variable $\gamma$. For the approximation bound to be conservative, it requires:
$\gamma\leq 2c+1.$ (8)
Given the right-skewed nature of $\gamma\geq 1$, we model $\gamma=1+X$, where $X\sim\text{Gamma}(\alpha,\beta)$. The probability that the approximation is valid is:
$P(\gamma\leq 2c+1)=P(X\leq 2c)=\frac{\gamma(\alpha,\beta\cdot 2c)}{\Gamma(\alpha)},$ (9)
where $\gamma(\alpha,z)=\int_{0}^{z}t^{\alpha-1}e^{-t}\,dt$ is the lower incomplete gamma function, and $\Gamma(\alpha)$ is the gamma function. The choice of $c$ trades off reliability and tightness:
  • If $c$ is small (e.g., $2c<\mathbb{E}[X]=\alpha/\beta$), the probability $P(X\leq 2c)$ is low, risking an invalid bound.
  • If $c$ is large, $P(X\leq 2c)\to 1$, but the bound $\beta\geq 1-\sqrt{cH_{q}}$ becomes looser.
Optimal performance requires balancing the reliability of the approximation (high $P(X\leq 2c)$) with the tightness of the bound (small $c$).
We analyze such a trade-off on Qwen2.5 on MT-Bench and QwQ-32B on AIME in Figure˜4. It shows that for most cases our estimated approximation bound works well (has a higher acceptance probability than the oracle bound when hyperparameter $c$ (i.e. $h$) is properly selected, set to 0.18 in the figure).
@@ 3 Experiments
Next, to verify the effectiveness of SVIP, we conduct experiments on both conventional long-form generation (Section 3.1) and reasoning with test-time scaling (Section 3.2). Since SVIP is completely training-free, we also apply it to other speculative decoding methods and demonstrate its flexibility (Section 3.1).
As baselines, we consider two widely adopted policies for draft length:
  1. 1.
    Constant: a constant draft length (set to 5 unless otherwise stated), which is commonly used in the literature
  2. 2.
@@
    Heuristic: the heuristics implemented in Hugging Face Transformers library (Wolf et al., 2019), where the draft length for the next draft iteration is increased by 2 if all draft tokens in the current iteration are accepted, and otherwise decreased by 1.
3.1 Results on Long-form Generation
Settings